OpenStudy (anonymous):

-4r - 11 = 4r + 21 please show work

OpenStudy (anonymous):

-4r-11=4r+21 -4r-4r=21+11 -8r=32 r=-32/8 r=-4

OpenStudy (anonymous):

Thx